# Hyperpolygonal arrangements

### v2

@@ 2.1. Hyperplane arrangements
A property for arrangements is said to be local if it is passed to every localization of the ambient arrangement. For instance, freeness, asphericity, and factoredness are all local (cf. [OT92, Thm. 4.37], [Pa93, Lem. 1.1], and the proof of [Ter92, Cor. 2.11], respectively). In contrast, this fails for formal arrangements of rank at least $4$ (cf. the example following [Y93, Def. 2.3]).
@@
An $\ell$-arrangement ${\mathscr{A}}$ with $\operatorname{rank}({\mathscr{A}})=r$ is called generic if every subarrangement ${\mathscr{B}}$ of ${\mathscr{A}}$ of cardinality $r$ is linearly independent and $|{\mathscr{A}}|>r$, [OT92, Def. 5.22].
@@ 2.3. Free arrangements
(i). Generic arrangements of rank greater than $2$ are not free, e.g., see [RT91, §4.4].
@@ 2.7. $K(\pi,1)$-arrangements
Let =${\mathscr{B}}=\{\ker x_{1},\ldots,\ker x_{n-1},\ker(x_{1}+\ldots+x_{n})\}\subseteq{\mathscr{H}}_{n}$. Then it is easy to see that successively all $H_{I}$ for $\varnothing\neq I\subseteq[n]$ belong to the line-losure ${\operatorname{lc}}({\mathscr{B}})$ of ${\mathscr{B}}$, as follows. For $i\in[n]$, set $H_{i}:=\ker x_{i}$, and for $\alpha:=x_{1}+\ldots+x_{n}$ set $H_{\alpha}:=\ker\alpha$. Then for $I=[n]\setminus\{i\}$, we have $H_{I}=\ker\left(\alpha-2x_{i}\right)\supset H_{\alpha}\cap H_{i}$. Thus for each $I$ of cardinality $n-1$, $H_{I}$ belongs to ${\operatorname{lc}}({\mathscr{B}})$. Now iterate the argument above for $I$ of successively smaller cardinality. Consequently, ${\operatorname{lc}}({\mathscr{B}})={\mathscr{H}}_{n}$. Since $\operatorname{rk}({\mathscr{H}}_{n})=n=|{\mathscr{B}}|=\operatorname{rk}({\mathscr{B}})$, it follows from Proposition 5.2 that ${\mathscr{H}}_{n}$ is combinatorially formal. ∎
@@ 6. Rank-generating functions of the poset of regions
Let $W$ be a Coxeter group again with reflection arrangement ${\mathscr{A}}={\mathscr{A}}(W)$, let $X$ be a member of the intersection lattice $L({\mathscr{A}})$, and consider the restricted reflection arrangement ${\mathscr{A}}^{X}$. In general, ${\mathscr{A}}^{X}$ is no longer a reflection arrangement. Nevertheless, thanks to [MR19, Thm. 1.3], there always exists a suitable base region $B$ of ${\mathscr{A}}^{X}$ in $X$ so that also $\zeta(P({\mathscr{A}}^{X},B);t)$ satisfies (6.1), provided $W$ is not of type $E_{8}$. In case $W$ is of type $E_{8}$, then $\zeta(P({\mathscr{A}}^{X},B);t)$ satisfies (6.1), provided $X$ has rank at most $3$ with only two exceptions or ${\mathscr{A}}^{X}\cong(E_{8},D_{4})$.
@@
It follows from Theorem 1.2(ii) and [JP95, Prop. 3.4, Thm. 6.1] that ${\mathscr{H}}_{n}$ satisfies (6.1) for $n\leq 3$, and for ${\mathscr{H}}_{4}$ this follows from Lemma 3.2 and (6.3). One can check that in contrast (6.1) fails for $n=5$.
Acknowledgments: We are grateful to T. Hoge and S. Wiesner for helpful discussions on material in this paper. The research of this work was supported in part by the DFG (Grant #RO 1072/25-1 (project number: 539865068) to G. Röhrle, as well as Grant MU #5286/1-1 (project number: 539874788) to P. Mücksch.

### v4

@@ 2.1. Hyperplane arrangements
An $\ell$-arrangement ${\mathscr{A}}$ with $\operatorname{rank}({\mathscr{A}})=r$ is called generic if every subarrangement ${\mathscr{B}}$ of ${\mathscr{A}}$ of cardinality $\ell$ is linearly independent, [OT92, Def. 5.22].
@@ 2.3. Free arrangements
(i). Generic arrangements are not free, e.g., see [RT91, §4.4].
@@ 2.7. $K(\pi,1)$-arrangements
Let =${\mathscr{B}}=\{\ker x_{1},\ldots,\ker x_{n-1},\ker(x_{1}+\ldots+x_{n})\}\subseteq{\mathscr{H}}_{n}$. Then it is easy to see that successively all $H_{I}$ for $\varnothing\neq I\subseteq[n]$ belong to the line-closure ${\operatorname{lc}}({\mathscr{B}})$ of ${\mathscr{B}}$, as follows. For $i\in[n]$, set $H_{i}:=\ker x_{i}$, and for $\alpha:=x_{1}+\ldots+x_{n}$ set $H_{\alpha}:=\ker\alpha$. Then for $I=[n]\setminus\{i\}$, we have $H_{I}=\ker\left(\alpha-2x_{i}\right)\supset H_{\alpha}\cap H_{i}$. Thus for each $I$ of cardinality $n-1$, $H_{I}$ belongs to ${\operatorname{lc}}({\mathscr{B}})$. Now iterate the argument above for $I$ of successively smaller cardinality. Consequently, ${\operatorname{lc}}({\mathscr{B}})={\mathscr{H}}_{n}$. Since $\operatorname{rk}({\mathscr{H}}_{n})=n=|{\mathscr{B}}|=\operatorname{rk}({\mathscr{B}})$, it follows from Proposition 5.2 that ${\mathscr{H}}_{n}$ is combinatorially formal. ∎
@@ 6. Rank-generating functions of the poset of regions
Let $W$ be a Coxeter group again with reflection arrangement ${\mathscr{A}}={\mathscr{A}}(W)$, let $X$ be a member of the intersection lattice $L({\mathscr{A}})$, and consider the restricted reflection arrangement ${\mathscr{A}}^{X}$. In general, ${\mathscr{A}}^{X}$ is no longer a reflection arrangement. It was shown in [MR19, Thm. 1.3] that there always exists a suitable base region $B$ of ${\mathscr{A}}^{X}$ in $X$ so that also $\zeta(P({\mathscr{A}}^{X},B);t)$ satisfies (6.1), with the exception of only three instances when $W$ is of type $E_{8}$.
@@
It follows from Theorem 1.2(ii) and [JP95, Prop. 3.4, Thm. 6.1] that ${\mathscr{H}}_{n}$ satisfies (6.1) for $n\leq 3$, and for ${\mathscr{H}}_{4}$ this follows from Lemma 3.2 and (6.3). One can check that in contrast (6.1) fails for ${\mathscr{H}}_{5}$.
Acknowledgments: We are grateful to T. Hoge and S. Wiesner for helpful discussions on material in this paper. The research of this work was supported in part by the DFG (Grant #RO 1072/25-1 (project number: 539865068) to G. Röhrle).